OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: y = x − 6 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why?

OpenStudy (bradely):

Line A: y = x − 6 Line B: y = 3x + 4 x-6 =3x+4 3x-x =-6-4 2x = -10 x =-5 y = x-6=-5-6 =-11 (-5,-11) Source: http://www.mathskey.com/question2answer/